OpenStudy (anonymous):

I need to make a Half-Life function out of this. F(x) = 37.23e^-0.55x y = 37.23e^-0.55x y/37.23 = e^(-0.55x) ln y/37.23 = ln e^(-0.55x) ln y – ln 37.23 = ln e^(-0.55x) (3.62) ln y = ln e^(-0.55x) + 3.62 ln y = -0.55x + 3.62

OpenStudy (anonymous):

@satellite73 hopefully you can help me.

OpenStudy (anonymous):

@hba, I am still looking for assistance! D:

OpenStudy (anonymous):

\[F(x) = 37.23e^{-0.55x}\]

OpenStudy (anonymous):

set \[.5=e^{-.55x}\] and solve for \(x\) to find the half life

OpenStudy (anonymous):

take two steps 1) take the log 2) divide by \(-.55\)

OpenStudy (anonymous):

\[e^{-.55x}=.5\] \[-.55x=\ln(.5)\] \[x=\frac{\ln(.5)}{-.55}\] then a calculator

OpenStudy (anonymous):

So I have to solve for X and fill in the blanks you left for me?

OpenStudy (anonymous):

half life is the time it takes to get half of what you started with the coefficient in front of the \(e\) is unimportant, since half of it will be, well, half in general if you have \[Pe^{kt}\] and you want the half life, half of \(P\) is \(\frac{P}{2}\) so you could start with \[Pe^{kt}=\frac{P}{2}\] but the first step is to divide by \(P\) and you get \[e^{kt}=\frac{1}{2}=.5\] so you might as well start with \[e^{kt}=.5\] go to \[kt=\ln(.5)\] and so \[t=\frac{\ln(.5)}{k}\]

OpenStudy (anonymous):

I am so lost, but I do appreciate the help haha.

OpenStudy (anonymous):

the problem with your solution above is that you kept the variable \(y\) but for the half life, \(y\) is not a variable, it is known to you namely, it is half of what you started with

OpenStudy (anonymous):

lets start from scratch because it is not too hard

OpenStudy (anonymous):

So we start with F(x) = 37.23e^-0.55x?

OpenStudy (anonymous):

you have \[F(x) = 37.23e^{-0.55x}\] as your original decay function

OpenStudy (anonymous):

now you want the half life i.e. how long before you have half of what you start with

OpenStudy (anonymous):

Couldn't I just divide it by 2, or do I need to have something else to go with it?

OpenStudy (anonymous):

half of \(37.25\) is \(\frac{37.25}{2}\) whatever that is

OpenStudy (anonymous):

so now we know what \(F(x)\) is, it is \(F(x)=\frac{37.25}{2}\) that is a number known to us

OpenStudy (anonymous):

It comes out to 18.615

OpenStudy (anonymous):

what we want to solve for is \(x\) that is, we want to solve \[\frac{37.25}{2}=37.25e^{-.55x}\]

OpenStudy (anonymous):

By the way, its 37.23! You confused me for a second.

OpenStudy (anonymous):

ok whatever, the point is that it does not matter what you start with

OpenStudy (anonymous):

So 37.23/2 = 18.6 which is equal to 37.23e^-.55x?

OpenStudy (anonymous):

because the first step in solving \[\frac{37.23}{2}=37.23e^{-.55x}\] is to divide both sides by \(37.23\)

OpenStudy (anonymous):

so the \(37.23\) is irrelevant to the problem, you should go right to \[\frac{1}{2}=e^{-.55x}\]

OpenStudy (anonymous):

make sure you understand that the initial amount, the number in front of the \(e\) is not important to the problem whatever you have there, half of it is half of it, so when you divide both sides by that number, you will get \(\frac{1}{2}\) no matter what number you start with

OpenStudy (anonymous):

if you have \[100e^{-.55x}\] then you would say half of 100 is 50 and get \[100e^{-.55x}=50\] which still comes down to solving \[e^{-.55x}=\frac{1}{2}\]

OpenStudy (anonymous):

Okay, so 1/2 = e^-.55x

OpenStudy (anonymous):

right now you are done in two steps take the log get \[\ln(.5)=-.55x\] and finally \[x=\frac{\ln(.5)}{-.55}\] it always works this way takes two steps only

OpenStudy (anonymous):

-1.243147181?

OpenStudy (anonymous):

the half life should be positive this is what i get http://www.wolframalpha.com/input/?i=-ln%28.5%29%2F.55

OpenStudy (anonymous):

Never knew about that site thanks for that, I think I just put it in incorrectly on my calculator xD

OpenStudy (anonymous):

wolf is the best way to cheat i have found tells you everything

OpenStudy (anonymous):

of course you have to know what to ask it here i solved directly without taking the log http://www.wolframalpha.com/input/?i=.5%3De^%28-.55x%29

OpenStudy (anonymous):

So that would be the final answer for that part?

OpenStudy (anonymous):

Once again @satellite73 I really appreciate the help thank you.